OpenStudy (anonymous):

Find the volume between z = square-root(x^2 + y^2) and z = (3/7) square-root(x^2 + y^2) and inside x^2 + y^2 + z^2 = 4 (a) Set up the integral to find the volume.

ganeshie8 (ganeshie8):

so what kind of surfaces do you have ?

OpenStudy (anonymous):

a cone ? @ganeshie8

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

so ok, it looks like a cone and the volume of the cone is pi r^2 (h/3)

OpenStudy (anonymous):

and they are asking to set the integral in spherical coordinates

ganeshie8 (ganeshie8):

Looks you have two cones and one sphere

OpenStudy (anonymous):

this is what know : \[\int\limits_{0}^{2\pi}\int\limits_{?}^{?}\int\limits_{0}^{2} f(\rho, \phi, \theta)\rho^2 \sin(\phi)d \rho d \phi d \theta \]

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

looks the radius is changing from 0 to 2 ?

ganeshie8 (ganeshie8):

Oh nice you have already figured out \(\rho \) and \(\theta\)

OpenStudy (anonymous):

i dont know phi

ganeshie8 (ganeshie8):

find the angle made by top cone with z axis

ganeshie8 (ganeshie8):

that gives you lower bound of \(\phi\)